# DyBBT: Dynamic Balance via Bandit inspired Targeting for Dialog Policy with Cognitive Dual-Systems

### v1

@@ A.1 Formalization of Cognitive State Space
  • User Uncertainty ($u_{t}=|S_{unconfirmed}|/|S_{relevant}|$) operationalizes the information gathering affordance. A high $u_{t}$ indicates ambiguity in the user’s goal, directly signaling the need for information seeking actions to reduce entropy, a well established principle in decision theory.
  • Slot Dependency ($\rho_{t}=\max_{u\in U}(\frac{1}{|F|}\sum_{f\in F}M(u,f))$) captures the structural affordance of the task environment, derived from a pre-computed slot co-occurrence matrix $M$ from the training corpus. A high $\rho_{t}$ suggests that the next piece of information is highly predictable given what is already known (e.g., requesting departure after knowing destination in a taxi domain), making targeted exploitation more efficient than random exploration. This dimension encodes the latent structure of the domain.
This design transforms the complex, unstructured exploration problem in the raw belief space into a more manageable one in a structured space where states with similar exploration needs are grouped together, as visualized in Figure 5.
@@ A.2 Regret Analysis Under Simplifying Assumptions
To provide theoretical intuition for our exploration principle, we present a regret analysis under a set of simplifying assumptions that capture the core structure that we aim to exploit. This analysis justifies the form of our exploration bonus and provides an upper bound on learning speed. We make the following assumptions to bridge the gap between bandit theory and the dialog POMDP. Our analysis is based on the Assumption 1 stated in Section 3.1.2, which posits Lipschitz smoothness of the reward function in the cognitive state space $\mathcal{C}$.
Assumption 2 (MDP over $\mathcal{C}$).
@@
$V^{\pi}(\mathbf{c})=\mathbb{E}\left[\sum_{k=0}^{H}\gamma^{k}\bar{r}(\mathbf{c}_{t+k},a_{t+k})\,\middle|\,\mathbf{c}_{t}=\mathbf{c},a_{t+k}\sim\pi(\cdot|\mathbf{c}_{t+k})\right].$
@@ A.2.1 Theoretical Intuition for Regret
Under Assumptions 1 and 2, if we perform optimistic exploration in the cognitive state space $\mathcal{C}$, prioritizing states with low visitation counts, we can derive an upper bound on the expected cumulative regret that scales sublinearly with time:
@@
where $R(T)=\sum_{t=1}^{T}[V^{*}(\mathbf{c}_{t})-V^{\pi_{t}}(\mathbf{c}_{t})]$ is the cumulative regret, and $\widetilde{\mathcal{O}}$ hides logarithmic factors. The notation $\lesssim$ indicates that this is a heuristic bound that captures the expected asymptotic scaling rather than a rigorous inequality. Here, $L_{r}$ is the Lipschitz constant from Assumption 1, bounding the reward’s sensitivity to changes in $\mathcal{C}$.
@@ A.2.2 Derivation Sketch
    Discretization Error: Due to Lipschitz continuity of $\bar{r}(\mathbf{c},a)$ (Assumption 1), the error introduced by discretization is bounded by $\mathcal{O}(L_{r}\epsilon T)$.
@@
This bound provides an intuitive justification for our exploration criterion (Eq. 1 in the main text). The term $\sqrt{\frac{\log T}{n_{t}(\mathbf{c}_{t})}}$ is a heuristic adaptation of the optimism principle, encouraging exploration of states with high uncertainty, inversely proportional to their visitation count. The empirical regret curve (Figure 7) shows sublinear growth, consistent with this theoretical intuition.
@@ A.3 Justification for the Meta-Controller Rule
$\text{Activate System 2 IF: }\left(n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}\right)\lor\left(p_{t}^{S1}<\kappa\right).$
@@
  • Model Imperfection: System 1, as a parameterized policy, may have inherent limitations and blind spots not captured by the visitation count.
@@
A low confidence score $p_{t}^{S1}$ is a proxy for these forms of uncertainty. This condition ensures robustness by invoking the powerful, knowledge rich System 2 when System 1 is uncertain, preventing catastrophic failures. The disjunctive ($\lor$) combination ensures System 2 is activated for either theoretical exploration or empirical robustness, making the overall system more adaptive and reliable than either condition alone, as evidenced by the ablation study (Table 5).
@@ A.4 Discussion and Limitations
Simplified Model: Assumption 2 reduces the POMDP to an MDP over $\mathcal{C}$, ignoring the challenges of belief state tracking and partial observability. This is a significant simplification. Our empirical results show that the algorithm performs well even when this assumption is not perfectly met, as the meta-controller’s confidence condition can mitigate some of these issues.
@@ B.1 Experimental Platform and Datasets
We evaluated DyBBT on two widely adopted benchmarks: the Microsoft dialog Challenge (MS dialog) (Li et al. (2018)) for single domain tasks, and the MultiWOZ 2.1 corpus (Budzianowski et al. (2018)) for multi domain tasks. Both datasets are converted into ConvLab-3’s unified format, ensuring consistency in ontology, state representation, and API interaction. Table 2 summarizes the key statistics of both datasets.
@@
The MultiWOZ 2.1 dataset is a large scale multi domain corpus spanning seven domains: Attraction, Hotel, Restaurant, Taxi, Train, Hospital, and Police. It includes 10,420 dialogs and 145,360 turns, with an average of 13.9 turns per dialog. The dataset is split into 8,420 dialogs for training, 1,000 for validation, and 1,000 for testing.
Both datasets provide annotated belief states, system dialog acts, and user goals, making them suitable for training and evaluating end-to-end dialog policies. The diversity in domain complexity, dialog length, and task structure across these datasets allows us to thoroughly assess the generalization capability of DyBBT in both single and multi domain settings.
To ensure reproducibility and enable fair comparison, we implement and evaluate our proposed DyBBT framework using ConvLab-3 Zhu et al. (2023), a flexible and unified toolkit for TODS. ConvLab-3 provides standardized data formats, integrated user simulators, and reinforcement learning utilities, facilitating consistent development and evaluation of dialog policies across multiple domains. All experiments are conducted using ConvLab-3’s builtin simulators and evaluation metrics, ensuring comparability across models and domains.
Table 2: Summary of dataset statistics for MS Dialog and MultiWOZ 2.1.
@@ B.2 Baselines Details
  • LLM_\_DP agents replace the dialog policy (DP) module of the TODS with GPT-4.0 (drawing on advances in LLM based multi turn dialog systems) to select appropriate actions and pass them to the natural language generation (NLG) module for response generation (Yi et al., 2024).
@@ B.3 Metrics Formula
This section provides the formal definitions of the evaluation metrics used for multi domain TODS evaluation, following the standard MultiWOZ evaluation protocol.
@@ B.3.1 Inform Success Rate
$\text{TP}=\sum_{d\in D}\sum_{s\in R_{d}}\mathbb{I}\left(\exists\,\text{inform}(d,s,v)\in S\land v\notin V_{\text{null}}\right)$ (4)
$\text{FP}=\sum_{d\in D}\sum_{s\notin R_{d}\cup I_{d}}\mathbb{I}\left(\exists\,\text{inform}(d,s,v)\in S\land v\notin V_{\text{null}}\right)$ (5)
$\text{FN}=\sum_{d\in D}\sum_{s\in R_{d}}\mathbb{I}\left(\nexists\,\text{inform}(d,s,v)\in S\lor v\in V_{\text{null}}\right)$ (6)
where $V_{\text{null}}=\{``",\text{``dont care"},\text{``not mentioned"}\}$ represents null values. The Inform Success Rate is then defined as:
@@ B.3.2 Book Success Rate
$\text{Book}_{d}=\frac{1}{|B_{d}|}\sum_{b\in B_{d}}\mathbb{I}\left(\text{book}(d,b,v)\in S\land v=v_{\text{goal}}\right)$ (8)
@@ B.4.1 System 1 Prompt
You are the fast, intuitive component (System 1) of a task oriented dialog system. Your task is to generate the next system action based solely on the current belief state. Do not reason step-by-step. Output your first, most intuitive response in the exact JSON format specified.
**Current Belief State:**
{belief_state}
**Available Actions:**
{available_actions}
Based on the above, output ONLY a valid JSON object with your predicted action and its confidence. Do not output any other text.
{"action": [["<act_type>", "<domain>", "<slot>"], ["<act_type>", "<domain>", "<slot>"], ...],"confidence": <confidence_score>}
@@ B.4.2 System 2 Prompt
You are the deliberative reasoner (System 2) of a task oriented dialog system. Your goal is to generate diverse, high quality action plans when the meta-controller detects a need for deeper reasoning, either due to unfamiliar cognitive states or low confidence from System 1.
**Current Belief State:**
{belief_state}
**Available Actions:**
{available_actions}
**Cognitive State Context:**
- dialog Progress: {d_t}
- User Uncertainty: {u_t}
- Slot Dependency: {p_t}
**Trigger Reason:** {trigger_reason}
**Reasoning Guidelines:**
1. **Leverage cognitive signals**:
- If progress is low, focus on information gathering.
- If uncertainty is high, prioritize clarifying or confirming actions.
- If slot dependency is high, leverage known slot relationships to guide next actions.
2. **Consider domain and slot dependencies**:
- E.g., taxi requires both destination and departure’; restaurant may require area’, food’, pricerange before booking.
3. **Generate 3 distinct strategies** that reflect different tactical approaches:
- One conservative (e.g., confirm before acting),
- One proactive (e.g., request multiple slots),
- One hybrid (e.g., inform then request).
4. **Evaluate each path** by estimating its likelihood of leading to task success.
**Output Format:** Strictly adhere to the following JSON schema:
"reasoning_paths": [
"sequence_id": 1,
"action_sequence": [
["action_type", "domain", "slot"],
"estimated_success_probability": 0.9
@@ B.4.3 LLM_DP Prompt
You must strictly execute the following commands:
1. Command execution requirements: when receiving a command, you must strictly follow the given instructions without performing any actions outside the scope of the command or generating any additional words.
2. Datasets and system roles: as the dialog policy component in a task oriented dialog system, you will make system decisions based on the MultiWOZ 2.1 dataset.
3. Processing user dialog state: you will receive a formatted user dialog state. This state will be used as a basis for decision making.
4. Generate system actions: based on the user dialog state {
user_action’: [["Inform", "Hotel", "Area", "east"], ["Inform", "Hotel", "Stars", "4"]],
system_action’: [],
belief_state’: {
police’: {’book’: {’booked’: []}, semi’: {}},
hotel’: {’book’: {’booked’: [], people’: ’’, day’: ’’, stay’: ’’},
semi’: {’name’: ’’, area’: east’, parking’: ’’, pricerange’: ’’, stars’: ’4’, internet’: ’’, type’: ’’}},
attraction’: {’book’: {’booked’: []}, semi’: {’type’: ’’, name’: ’’, area’: ’’}},
restaurant’: {’book’: {’booked’: [], people’: ’’, day’: ’’, time’: ’’},
semi’: {’food’: ’’, pricerange’: ’’, name’: ’’, area’: ’’}},
hospital’: {’book’: {’booked’: []}, semi’: {’department’: ’’}},
taxi’: {’book’: {’booked’: []},
semi’: {’leaveAt’: ’’, destination’: ’’, departure’: ’’, arriveBy’: ’’}},
train’: {’book’: {’booked’: [], people’: ’’},
semi’: {’leaveAt’: ’’, destination’: ’’, day’: ’’, arriveBy’: ’’, departure’: ’’}}
request_state’: {},
terminated’: False,
history’: []
}, you need to generate system actions. These actions should be provided in the following format: [["ActionType", "Domain", "Slot", "Value"]] where ActionType denotes the type of action (e.g. Request, Inform, Confirm, etc.), Domain specifies the associated domain (e.g. restaurant, taxi, hotel, etc.), Slot is the specific information slot associated with the action (e.g. name, area, type, etc.), and Value is the corresponding value or an empty string.
@@ B.5 Implementation Details
The DyBBT framework was implemented within the Convlab-3 dialog system environment (Zhu et al. (2023)), leveraging its modular architecture for efficient dialog policy optimization. We employed RuleDST for system dialog state tracking and RulePolicy for user policy simulation, eliminating the need for natural language understanding (NLU) and natural language generation (NLG) modules. This design choice significantly enhances training efficiency by reducing computational overhead and isolating the impact of language processing components from policy learning performance. The dialog environment was configured with a maximum turn limit of 30 for single domain and 40 for multi domain (the same as EIERL) interactions per episode, with the cognitive state space $\mathcal{C}$ computed in real-time during dialog execution using dimensions including dialog progress ($d_{t}$), user uncertainty ($u_{t}$), and slot dependency ($\rho_{t}$) extracted from the belief state representation provided by RuleDST.
@@
All experiments were conducted on NVIDIA 5090 GPUs with 32GB memory. System 1 was SFT using the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and further optimized via PPO, employing a clipping parameter $\epsilon=0.2$ and GAE with $\lambda=0.95$. The meta-controller employs a dual-threshold mechanism for System 2 invocation, with $kappa=0.7$ and $\tau=1.0$, values selected via grid search over development sets as they maximize both performance and robustness across domains. These thresholds operate on a discretized 5 bins cognitive state space, which balances expressiveness and generalization, as validated in Section D.4.
We maintained a replay buffer with a capacity of 10,000 transitions, using a batch size of 32 for training. A separate knowledge distillation buffer was managed under a FIFO replacement policy with a fixed capacity. To ensure reproducibility, all experiments were run with five fixed random seeds (9841, 35741, 91324, 8134, 13924), consistent with the EIERL baseline Zhao et al. (2025). All hyperparameters were selected through grid search on a validation subset of the MultiWOZ data.
Training was conducted for 500 epochs on single domain tasks and 10,000 epochs on multi domain tasks, incorporating early stopping with a patience of 3 epochs based on validation performance. This protocol aligns with the EIERL setup for fair comparison.
@@ B.5.1 Slot Co-occurrence Matrix Construction
The slot dependency dimension $\rho_{t}$ in the cognitive state space $\mathcal{C}$ is derived from a co-occurrence matrix $M$ that captures statistical relationships between dialog slots across the Microsoft dialog Challenge (Li et al. (2018)) and MultiWOZ (Budzianowski et al. (2018)) dataset. This matrix quantifies the conditional probability that slot $j$ appears given the presence of slot $i$, providing a principled measure of semantic relatedness between dialog concepts.
@@
B.5.2 Knowledge Distillation Buffer Management

### v2

@@ A.1 Formalization of Cognitive State Space
  • User Uncertainty operationalizes the information gathering affordance.
    $u_{t}=|S_{unconfirmed}|/|S_{relevant}|$
    A high $u_{t}$ indicates ambiguity in the user’s goal, directly signaling the need for information seeking actions to reduce entropy, a well established principle in decision theory.
  • Slot Dependency captures the structural affordance of the task environment, derived from a pre-computed slot co-occurrence matrix $M$ from the training corpus.
    $\rho_{t}=\max_{u\in U}(\frac{1}{|F|}\sum_{f\in F}M(u,f))$
    A high $\rho_{t}$ suggests that the next piece of information is highly predictable given what is already known (e.g., requesting departure after knowing destination in a taxi domain), making targeted exploitation more efficient than random exploration. This dimension encodes the latent structure of the domain.
This design transforms the complex, unstructured exploration problem in the raw belief space into a more manageable one in a structured space where states with similar exploration needs are grouped together, as visualized in Fig. 5.
@@ A.2 Regret Analysis Under Simplifying Assumptions
To provide theoretical intuition for our exploration principle, we present a regret analysis under a set of simplifying assumptions that capture the core structure that we aim to exploit. This analysis justifies the form of our exploration bonus and provides an upper bound on learning speed. We make the following assumptions to bridge the gap between bandit theory and the dialog POMDP. Our analysis is based on the Assumption 3.1 stated in Sec. 3.1.2, which posits Lipschitz smoothness of the reward function in the cognitive state space $\mathcal{C}$.
Assumption A.1 (MDP over $\mathcal{C}$).
@@
$V^{\pi}(\mathbf{c})=\mathbb{E}\!\left[\sum_{k=0}^{H}\gamma^{k}\,\bar{r}(\mathbf{c}_{t+k},a_{t+k})\;\middle|\;\mathbf{c}_{t}=\mathbf{c},\right.\\ \left.a_{t+k}\sim\pi(\cdot|\mathbf{c}_{t+k})\right].$ (4)
@@ A.2.1 Theoretical Intuition for Regret
Under Assumptions 3.1 and A.1, if we perform optimistic exploration in the cognitive state space $\mathcal{C}$, prioritizing states with low visitation counts, we can derive an upper bound on the expected cumulative regret that scales sublinearly with time:
@@
where $R(T)=\sum_{t=1}^{T}[V^{*}(\mathbf{c}_{t})-V^{\pi_{t}}(\mathbf{c}_{t})]$ is the cumulative regret, and $\widetilde{\mathcal{O}}$ hides logarithmic factors. The notation $\lesssim$ indicates that this is a heuristic bound that captures the expected asymptotic scaling rather than a rigorous inequality. Here, $L_{r}$ is the Lipschitz constant from Assumption 3.1, bounding the reward’s sensitivity to changes in $\mathcal{C}$.
@@ A.2.2 Derivation Sketch
    Discretization Error: Due to Lipschitz continuity of $\bar{r}(\mathbf{c},a)$ (Assumption 3.1), the error introduced by discretization is bounded by $\mathcal{O}(L_{r}\epsilon T)$.
@@
This bound provides an intuitive justification for our exploration criterion (Eq. 2 in the main text). The term $\sqrt{\frac{\log T}{n_{t}(\mathbf{c}_{t})}}$ is a heuristic adaptation of the optimism principle, encouraging exploration of states with high uncertainty, inversely proportional to their visitation count. The empirical regret curve (Fig. 7) shows sublinear growth, consistent with this theoretical intuition.
@@ A.3 Justification for the Meta-Controller Rule
Activate S2 IF:
$\left(n_{t}(\mathbf{c}_{t})<\tau\sqrt{\log T}\right)\lor\left(p_{t}^{S1}<\kappa\right).$ (6)
@@
  • Model Imperfection: S1, as a parameterized policy, may have inherent limitations and blind spots not captured by the visitation count.
@@
A low confidence score $p_{t}^{S1}$ is a proxy for these forms of uncertainty. This condition ensures robustness by invoking the powerful, knowledge rich S2 when S1 is uncertain, preventing catastrophic failures. The disjunctive ($\lor$) combination ensures System 2 is activated for either theoretical exploration or empirical robustness, making the overall system more adaptive and reliable than either condition alone, as evidenced by the ablation study (Table 2).
@@ A.4 Discussion and Limitations
Simplified Model: Assumption A.1 reduces the POMDP to an MDP over $\mathcal{C}$, ignoring the challenges of belief state tracking and partial observability. This is a significant simplification. Our empirical results show that the algorithm performs well even when this assumption is not perfectly met, as the meta-controller’s confidence condition can mitigate some of these issues.
@@ B.1 Experimental Platform and Datasets
We evaluated DyBBT on two widely adopted benchmarks: the Microsoft dialog Challenge (MS dialog) (Li et al. (2018)) for single-domain tasks, and the MultiWOZ 2.1 corpus (Eric et al. (2020)) for multi-domain tasks. Both datasets are converted into ConvLab-3’s unified format, ensuring consistency in ontology, state representation, and API interaction. Table 3 summarizes the key statistics of both datasets.
@@
The MultiWOZ 2.1 dataset is a large scale multi-domain corpus spanning seven domains: Attraction, Hotel, Restaurant, Taxi, Train, Hospital, and Police. It includes 10,420 dialogs and 145,360 turns, with an average of 13.9 turns per dialog. The dataset is split into 8,420 dialogs for training, 1,000 for validation, and 1,000 for testing.
Both datasets provide annotated belief states, system dialog acts, and user goals, making them suitable for training and evaluating end-to-end dialog policies. The diversity in domain complexity, dialog length, and task structure across these datasets allows us to thoroughly assess the generalization capability of DyBBT in both single and multi-domain settings.
To ensure reproducibility and enable fair comparison, we implement and evaluate our proposed DyBBT framework using ConvLab-3 (Zhu et al., 2023), a flexible and unified toolkit for TODS. ConvLab-3 provides standardized data formats, integrated user simulators, and reinforcement learning utilities, facilitating consistent development and evaluation of dialog policies across multiple domains. All experiments are conducted using ConvLab-3’s builtin simulators and evaluation metrics, ensuring comparability across models and domains.
@@
Table 3: Summary of dataset statistics for MS Dialog and MultiWOZ 2.1.
@@ B.2 Baselines Details
  • LLM_\_DP agents replace the dialog policy (DP) module of the TODS with GPT-4.0 (drawing on advances in LLM based multi turn dialog systems) to select appropriate actions and pass them to the natural language generation (NLG) module for response generation (Yi et al., 2025).
@@ B.3 Metrics Formula
This section provides the formal definitions of the evaluation metrics used for multi-domain TODS evaluation, following the standard MultiWOZ evaluation protocol.
@@ B.3.1 Inform Success Rate
$\text{TP}=\sum_{d\in D}\sum_{s\in R_{d}}\mathbb{I}\left(\exists\,\text{inform}(d,s,v)\right.\\ \left.\in S\land v\notin V_{\text{null}}\right)$ (7)
$\text{FP}=\sum_{d\in D}\sum_{s\notin R_{d}\cup I_{d}}\mathbb{I}\left(\exists\,\text{inform}(d,s,v)\right.\\ \left.\in S\land v\notin V_{\text{null}}\right)$ (8)
$\text{FN}=\sum_{d\in D}\sum_{s\in R_{d}}\mathbb{I}\left(\nexists\,\text{inform}(d,s,v)\right.\\ \left.\in S\lor v\in V_{\text{null}}\right)$ (9)
where $V_{\text{null}}=\{``^{\prime\prime},\text{``dont care''},\text{``not mentioned''}\}$ represents null values. The Inform Success Rate is then defined as:
@@ B.3.2 Book Success Rate
$\text{Book}_{d}=\frac{1}{|B_{d}|}\sum_{b\in B_{d}}\mathbb{I}\left(\text{book}(d,b,v)\right.\\ \left.\in S\land v=v_{\text{goal}}\right)$ (11)
@@ B.5 Implementation Details
The DyBBT framework was implemented within the Convlab-3 dialog system with Python 3.10 environment (Zhu et al. (2023)), leveraging its modular architecture for efficient dialog policy optimization. We employed RuleDST for system dialog state tracking and RulePolicy for user policy simulation, eliminating the need for natural language understanding (NLU) and natural language generation (NLG) modules. This design choice significantly enhances training efficiency by reducing computational overhead and isolating the impact of language processing components from policy learning performance. The dialog environment was configured with a maximum turn limit of 30 for single-domain and 40 for multi-domain (the same as EIERL) interactions per episode, with the cognitive state space $\mathcal{C}$ computed in real-time during dialog execution using dimensions including dialog progress ($d_{t}$), user uncertainty ($u_{t}$), and slot dependency ($\rho_{t}$) extracted from the belief state representation provided by RuleDST.
@@
All experiments were conducted on NVIDIA 5090 GPUs with 32GB memory. System 1 was SFT using the AdamW optimizer with a learning rate of $1\times 10^{-4}$ and further optimized via PPO, employing a clipping parameter $\epsilon=0.2$ and GAE with $\lambda=0.95$. The meta-controller employs a dual-threshold mechanism for System 2 invocation, with $kappa=0.7$ and $\tau=1.0$, values selected via grid search over development sets as they maximize both performance and robustness across domains. These thresholds operate on a discretized 5 bins cognitive state space, which balances expressiveness and generalization, as validated in Sec. E.5.
We maintained a replay buffer with a capacity of 10,000 transitions, using a batch size of 32 for training. A separate knowledge distillation buffer was managed under a FIFO replacement policy with a fixed capacity. To ensure reproducibility, all experiments were run with five fixed random seeds (9841, 35741, 91324, 8134, 13924), consistent with the EIERL baseline (Zhao et al., 2025). All hyperparameters were selected through grid search on a validation subset of the MultiWOZ data.
Training was conducted for 500 epochs on single-domain tasks and 10,000 epochs on multi-domain tasks, incorporating early stopping with a patience of 3 epochs based on validation performance. This protocol aligns with the EIERL setup for fair comparison.
@@ B.5.1 Slot Co-occurrence Matrix Construction
The slot dependency dimension $\rho_{t}$ in the cognitive state space $\mathcal{C}$ is derived from a co-occurrence matrix $M$ that captures statistical relationships between dialog slots across the Microsoft dialog Challenge (Li et al. (2018)) and MultiWOZ (Eric et al. (2020)) dataset. This matrix quantifies the conditional probability that slot $j$ appears given the presence of slot $i$, providing a principled measure of semantic relatedness between dialog concepts.
@@
B.5.2 Training Details For System 1
To train System 1 for accurate action prediction and confidence estimation, we employ a two-stage training methodology comprising supervised fine-tuning (SFT) followed by reinforcement learning. This approach utilizes dialog sequences from the MultiWOZ and MS Diag dataset to develop a robust policy model capable of rapid decision making with calibrated confidence scores.
Stage 1: Supervised Fine-tuning with Data Augmentation
We first construct a training corpus of 10,000 single turn dialog samples through systematic data augmentation. For each dialog turn, we extract the belief state $\mathbf{s}_{t}$, available action set $\mathcal{SA}$, and ground truth system actions $a_{t}^{*}$. The initial confidence score $p_{t}^{S1}$ is sampled from $\mathcal{U}(0.95,1.0)$.
The augmentation process introduces controlled perturbations to simulate prediction uncertainty. For each ground truth action sequence $a_{t}^{*}$, we apply three modification operations with specified probabilities: 20% action addition by sampling new actions from $\mathcal{SA}$; 60% action modification through substitution with random actions from $\mathcal{SA}$; and 20% action deletion while ensuring the augmented sequence $a_{t}^{\prime}$ maintains at least one action. These operations are applied sequentially in random order to each sample (Kadavath et al., 2022; Lin et al., 2022; Yin et al., 2023). The confidence score is adjusted proportionally to the modification intensity:
$p_{t}^{S1}\leftarrow p_{t}^{S1}\cdot\left(1-\frac{n_{\text{mod}}}{n}\right),$
where $n$ denotes the original action sequence length and $n_{\text{mod}}$ represents the number of modified actions. This procedure generates a dataset with confidence scores approximately uniformly distributed in $[0,1]$.
For SFT training, the model takes $\mathbf{s}_{t}$ and $\mathcal{SA}$ as inputs and produces both action sequence $a_{t}^{S1}$ and confidence score $p_{t}^{S1}$ as outputs. The composite loss function integrates action prediction and confidence estimation:
$\mathcal{L}=\lambda\mathcal{L}_{\text{a}}+(1-\lambda)\mathcal{L}_{\text{p}},$
where $\lambda=0.7$. The action loss $\mathcal{L}_{\text{a}}$ employs cross-entropy to measure divergence between predicted and augmented actions:
$\mathcal{L}_{\text{a}}=-\sum_{i}\log P(a_{t}^{S1}=a_{t}^{\prime}\mid\mathbf{s}_{t},\mathcal{SA}),$
while the confidence loss $\mathcal{L}_{\text{p}}$ utilizes mean squared error:
$\mathcal{L}_{\text{p}}=\left(p_{t}^{S1}-p_{t}^{\text{target}}\right)^{2}.$
Stage 2: Reinforcement Learning with PPO
The second stage employs PPO to optimize dialog level performance metrics using the complete MultiWOZ dataset. The reward function $R$ combines multiple objectives:
$R=R_{\text{success}}+R_{\text{efficiency}}+R_{\text{penalty}},$
where $R_{\text{success}}=+2t$ for successful dialogs and t-t for failures ($t$ denotes the max turn number), $R_{\text{efficiency}}=-1$ per dialog turn to encourage conciseness, and $R_{\text{penalty}}$ captures additional constraints.
This two-stage approach enables System 1 to initially learn accurate action confidence mappings through supervised learning, then refine its policy for improved task completion efficiency and success rates via reinforcement learning.
B.5.3 Knowledge Distillation Buffer Management